OpenStudy (anonymous):

What is the solution to the equation log 2 x – log 2 4 = 2

OpenStudy (campbell_st):

its really \[\log_{2} 4 = \log_{2} 2^2 = 2\] so the question becomes \[\log_{2} x -2 = 2\] or \[\log_{2}x = 4 \] so \[x = 2^4\]

OpenStudy (anonymous):

omg thank you i've been trying to do this question for the longets.

OpenStudy (campbell_st):

hope it helps

OpenStudy (zzr0ck3r):

you have been trying to do this question for the logest?

OpenStudy (zzr0ck3r):

ha, ill be at roccos comedy shack on tues and thurs.